# Nuclear Toeplitz operators between Fock spaces

### v1

@@ 3. Nuclear class Toeplitz operator
Let $1\leq q\leq p<\infty,1\leq s\leq\infty$, and let $\mu\geq 0$ satisfying (1.2), then the following statements are equivalent
  1. (1)
    $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$;
  2. (2)
    $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{s})$;
  3. (3)
    $\mu(\mathbb{C})<\infty$.
Moreover,
$\|T_{\mu}\|_{\mathcal{N}}\simeq\mu(\mathbb{C}).$ (3.1)
By a direct calculation, we see that $\widetilde{T_{\mu}}(z)=\widetilde{\mu}(z)$. Since $\mu\geq 0$, we can use Fubini’s Theorem to obtain
$\displaystyle\|\widetilde{\mu}\|_{L^{1}}$ $\displaystyle=\frac{\alpha}{\pi}\int_{\mathbb{C}}\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)dA(z)$
$\displaystyle=\frac{\alpha}{\pi}\int_{\mathbb{C}}\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}dA(z)d\mu(w)$
$\displaystyle=\mu(\mathbb{C}).$
$(1)\Rightarrow(3)$: If $1\leq q\leq p<\infty$, by Theorem 3.2 (3), we have $\widetilde{T_{\mu}}(z)=\widetilde{\mu}(z)\in L^{1}(\mathbb{C},dA)$, which implies $\mu(\mathbb{C})<\infty$. $(3)\Rightarrow(1)$ follows from Corollary 3.6. The proof of $(2)\Leftrightarrow(3)$ is similar.
Estimate (3.1) follows from the equation $\|\widetilde{\mu}\|_{L^{1}}=\mu(\mathbb{C})$, Theorem 3.2 and Theorem 3.5. ∎
Now, we present an interesting result. This conclusion is not immediately obvious, but it follows directly from our previous results.
Corollary 3.8.
Let $f\in L^{1}(\mathbb{C})$, then $\widetilde{f}\in\cap_{r\geq 1}L^{r}(\mathbb{C})$. Moreover, for any $r\geq 1$, we have $\|\widetilde{f}\|_{L^{r}}\lesssim\|f\|_{L^{1}}$.
It follows from Theorem 3.5, we have $T_{f}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$, then by Theorem 3.2, we have $\widetilde{f}=\widetilde{T_{f}}\in L^{r}(\mathbb{C})$ for all $\frac{1}{2}\leq r\leq\infty$. Moreover, applying Theorem 3.2 and Theorem 3.5, we obtain
$\|\widetilde{f}\|_{L^{r}}\lesssim\|T_{f}\|_{\mathcal{N}}\lesssim\|f\|_{L^{1}}.$
Let $\overline{F}_{\alpha}^{p}=\{\overline{f}:f\in F_{\alpha}^{p}\}$, suppose $\mu$ is a complex Borel measure and satisfies condition (1.1), we can define small Hankel operator on $K$ by
$h_{\mu}(f)(z)=\frac{\alpha}{\pi}\int_{\mathbb{C}}K_{\alpha}(w,z)f(w)\mathrm{e}^{-\alpha|w|^{2}}\,d\mu(w),\quad z\in\mathbb{C}.$
Theorem 3.9.
Let $\mu$ be a complex Borel measure on $\mathbb{C}$, and $|\mu|(\mathbb{C})<\infty$. Then $h_{\mu}\in\mathcal{N}(F_{\alpha}^{p},\overline{F}_{\alpha}^{q})(\mathcal{N}(f_{\alpha}^{\infty},\overline{F}_{\alpha}^{q}))$, where $1\leq p<\infty,1\leq q\leq\infty$. Moreover, we have
$\|h_{\mu}\|_{\mathcal{N}}\lesssim|\mu|(\mathbb{C}).$
Following the same method as in the proof of Theorem 3.5, we define
$h_{r}=\sum_{k=1}^{\infty}\mu(S_{r}+a_{k})\,k_{a_{k}}\otimes\overline{k_{a_{k}}}.$
Taking $r_{n}=\frac{1}{2^{n}}$, we have $h_{r_{n}}\to h_{\mu}$ in the nuclear norm $\|\cdot\|_{\mathcal{N}}$, which establishes the desired result.
Definition 3.10.
Let $X$ and $Y$ be Banach spaces. A linear operator $T\colon X\to Y$ is called absolutely summing provided there exists a constant $C>0$ such that for any finite family $\{x_{j}\}_{j=1}^{n}\subset X$,
$\sum_{j=1}^{n}\|Tx_{j}\|\;\leq\;C\sup_{\|x^{\prime}\|\leq 1}\sum_{j=1}^{n}|x^{\prime}(x_{j})|\;=\;C\sup_{|\varepsilon_{j}|=1}\Big\|\sum_{j=1}^{n}\varepsilon_{j}x_{j}\Big\|.$
Every nuclear operator is absolutely summing. In the special cases $X=c_{0}$ or $X=\ell_{\infty}$, the two classes coincide [8, Lemma 2.1].
Corollary 3.11.
Let $1\leq s\leq\infty$, and let $\mu\geq 0$ satisfying (1.2), then $T_{\mu}:f_{\alpha}^{\infty}\rightarrow F_{\alpha}^{s}$ is absolutely summing if and only if $\mu(\mathbb{C})<\infty$.
It is known that $f_{\alpha}^{\infty}$ is isomorphic to $c_{0}$(see [15]). Then the conclusion follows from Theorem 3.7. ∎
Remark 3.12.
Suppose $1\leq p<q,$ and $\mu\geq 0$. If $\mu(\mathbb{C})<\infty$, by Corollary 3.6, we have $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$. However, we cannot use Theorem 3.2 to prove the equivalence between $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $\mu(\mathbb{C})<\infty$, since in this case $(\frac{1}{p^{\prime}}+\frac{1}{q})^{-1}>1$. In fact, we will show that there exists $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ such that $\widetilde{T}(z)\notin L^{1}$.
Theorem 3.13.
Let $1\leq p<\infty$, and let $g$ is a entire function. Then
@@
Then we have $\varphi_{g_{n}}\in(F_{\alpha}^{p})^{\prime}$. Fix $f\in F_{\alpha}^{p}$, since $|g_{n}(z)|\leq|g(z)|$, we can use dominated convergence theorem to obtain
@@
Hence $\sup_{n}|\varphi_{g_{n}}(f)|<\infty$. The Principle of Uniform Boundedness yields
@@
for all $f\in F_{\alpha}^{p}$. It implies that $\varphi_{g}\in(F_{\alpha}^{p})^{\prime}$. Since we can identify the dual space $(F_{\alpha}^{p})^{\prime}$ with $F_{\alpha}^{p^{\prime}}$ through the pairing $\langle\cdot,\cdot\rangle_{F_{\alpha}^{2}}$, we can find $g^{\prime}\in F_{\alpha}^{p^{\prime}}$ such that
@@
The dominated convergence theorem then gives
$\displaystyle\varphi_{g-g^{\prime}}(z^{n})$ $\displaystyle=\int_{\mathbb{C}}z^{n}\overline{(g-g^{\prime})(z)}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=lim_{R\rightarrow{\infty}}\int_{|z|<R}z^{n}\overline{(g-g^{\prime})(z)}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=lim_{R\rightarrow{\infty}}\overline{a_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z).$
$\displaystyle=\overline{a_{n}}\int_{\mathbb{C}}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=0.$
@@
The other direction follows from Hölder’s inequality. ∎
@@
It is well-known that for $1\leq p<q\leq\infty$, we have the (proper) inclusion $F_{\alpha}^{p}\subset F_{\alpha}^{q}$. Theorem 3.13 tells us: We can find $f\in F_{\alpha}^{p^{\prime}}$ and $g\in F_{\alpha}^{q}$ such that
@@
The membership of $f$ in $F_{\alpha}^{p^{\prime}}$ is verified by computing two sums. For $1<p^{\prime}\leqslant 2$,
@@
$\sum_{n=1}^{\infty}\left|a_{n}\right|^{p}\left(\frac{n!}{\alpha^{n}}\right)^{\frac{p}{2}}n^{\frac{p}{4}-\frac{1}{2}}=\sum_{k=1}^{\infty}b^{pk}n_{k}^{\frac{p}{4}-\frac{1}{2}+\frac{p}{4}-\frac{p}{2p^{\prime}}}\epsilon_{n_{k}}^{p}\leq\sum_{k=1}^{\infty}\left(\frac{b^{p}}{2^{p}}\right)^{k}<\infty.$
The convergence of these series, together with Theorems 1 and 4 in [18], implies $f\in F_{\alpha}^{p^{\prime}}.$
Note that the coefficients satisfy
$a_{n}=\begin{cases}b^{k}\sqrt{\frac{\alpha^{n_{k}}}{(n_{k})!}}\,n_{k}^{\frac{1}{4}-\frac{1}{2q^{\prime}}}\,,&\text{if }n=n_{k},\\ 0,&\text{otherwise}.\end{cases}$
Consequently,
$a_{n}\neq o\left(\sqrt{\frac{\alpha^{n}}{n!}}n^{\frac{1}{4}-\frac{1}{2q^{\prime}}}\right)\quad(n\rightarrow\infty),$
which, by Corollary 5 of [18], means $f\in F_{\alpha}^{p^{\prime}}\setminus F_{\alpha}^{q^{\prime}}$.
@@
$b_{m}=\begin{cases}b^{k}\sqrt{\frac{\alpha^{n_{k}}}{(n_{k})!}}\,n_{k}^{\frac{1}{4}-\frac{1}{2q}}\,\epsilon_{n_{k}},&\text{if }m=n_{k},\\ 0,&\text{otherwise},\end{cases}$
@@
If $\int_{\mathbb{C}}|f(z){g(z)}|e^{-\alpha|z|^{2}}dA(z)<\infty$, Applying the dominated convergence theorem, we obtain
$\displaystyle\int_{\mathbb{C}}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}dA(z)$ $\displaystyle=lim_{R\rightarrow{\infty}}\int_{|z|<R}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=\lim_{R\rightarrow{\infty}}\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}a_{n}\overline{b_{m}}\int_{|z|<R}z^{n}\overline{z}^{m}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=\lim_{R\rightarrow{\infty}}\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z).$
From the construction above, we have $a_{n}\overline{b_{n}}\geq 0$ and $\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$ is monotonically increasing in $R$. The monotone convergence theorem yields:
$\displaystyle\int_{\mathbb{C}}f(z)\overline{g(z)}e^{\alpha|z|^{2}}dA(z)$ $\displaystyle=lim_{R\rightarrow{\infty}}\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{\mathbb{C}}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=\frac{\pi}{\alpha}\sum_{k=1}^{\infty}b^{2k}\epsilon_{n_{k}}$
$\displaystyle\geq\frac{\pi}{\alpha}\sum_{k=1}^{\infty}(\frac{b^{2}}{3})^{k}$
$\displaystyle=\infty.$
@@
However, for $1\leq p<q\leq\infty$ and $\mu\geq 0$, it remains unknown whether $\mu(\mathbb{C})<\infty$ is necessary for $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$. Finally, we conclude this section with an open problem.
Question 3.15.
@@
4. Trace formula and density theorem
@@
$Tr(T)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T}(z)dA(z).$
Let $T=\sum_{j=1}^{\infty}f_{j}\otimes g_{j}$, where $\left(g_{j}\right)\subset F_{\alpha}^{p},\left(f_{j}\right)\subset F_{\alpha}^{p^{\prime}}$ be a nuclear representation of $T$. Then we have $Tr(T)=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle_{F_{\alpha}^{2}}$. And the trace is independent of the representation choice. Then we have:
$\displaystyle\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T}(z)dA(z)$ $\displaystyle=\frac{\alpha}{\pi}\int_{\mathbb{C}}\langle Tk_{z},k_{z}\rangle_{F_{\alpha}^{2}}dA(z)$
$\displaystyle=\frac{\alpha}{\pi}\int_{\mathbb{C}}\sum_{j=1}^{\infty}\langle K_{z},f_{j}\rangle_{F_{\alpha}^{2}}\langle g_{j},K_{z}\rangle_{F_{\alpha}^{2}}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=\sum_{j=1}^{\infty}\int_{\mathbb{C}}\overline{f_{j}(z)}g_{j}(z)d\lambda_{\alpha}(z)$
$\displaystyle=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle_{F_{\alpha}^{2}}$
$\displaystyle=Tr(T).$
Proposition 4.2.
Suppose $\mu\geq 0$ and $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p})$, where $1<p<\infty$, then $Tr(T_{\mu})=\frac{\alpha}{\pi}\mu(\mathbb{C})$.
Let $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p})$, then we have $\mu(\mathbb{C})<\infty$ . By Fubini’s theorem, we obtain:
$Tr(T_{\mu})$ $=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{\mu}(z)dA(z)$
$=(\frac{\alpha}{\pi})^{2}\int_{\mathbb{C}}dA(z)\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)$
$=(\frac{\alpha}{\pi})^{2}\int_{\mathbb{C}}d\mu(w)\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}dA(z)$
$=\frac{\alpha}{\pi}\mu(\mathbb{C}).$
We now prove a useful result. The technical hypotheses in its statement are needed to justify the application of Fubini’s theorem. Notice that the case $p=q=2$ of the following theorem appears in [3, Theorem 6], but we adopt a different approach here.
@@
$Tr(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\widetilde{S}(z)dA(z).$
Let $f\in F_{\alpha}^{p}$, we can find a sequence $f_{n}\in K$ such that $f_{n}$ converges to $f$ with $\|\|_{\alpha,p}$. The set $K$ is defined as (1.3). By assumption (2), we have $T_{\varphi}f_{n}\stackrel{{\scriptstyle\|\cdot\|_{\alpha,q}}}{{\longrightarrow}}T_{\varphi}f$, hence $T_{\varphi}f_{n}(z)\rightarrow T_{\varphi}f(z)$ for every $z\in\mathbb{C}$.
@@
$\displaystyle|\int_{\mathbb{C}}(f_{n}(w)-f(w))\varphi(w)K(z,w)d\lambda_{\alpha}(w)|$
$\displaystyle\leq\frac{\alpha}{\pi}\int_{\mathbb{C}}|(f_{n}(w)-f(w))e^{-\frac{\alpha}{2}|z|^{2}}||\varphi(w)K(z,w)e^{-\frac{\alpha}{2}|z|^{2}}|dA(w)$
$\displaystyle\lesssim\|f_{n}-f\|_{\alpha,p}\|\varphi K_{z}\|_{\alpha,p^{\prime}}\rightarrow 0(n\rightarrow\infty).$
@@
$Tr(T_{\varphi}S)$ $=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T_{\varphi}S}(z)dA(z)$ (4.1)
@@
$\displaystyle\int_{\mathbb{C}}K_{u}(z)SK_{z}(u)d\lambda_{\alpha}(z)$ $\displaystyle=\int_{\mathbb{C}}K_{u}(z)\langle SK_{z},K_{u}\rangle_{F_{\alpha}^{2}}d\lambda_{\alpha}(z)$
$\displaystyle=\int_{\mathbb{C}}K_{u}(z)\langle K_{z},S^{\prime}K_{u}\rangle_{F_{\alpha}^{2}}d\lambda_{\alpha}(z)$
$\displaystyle=\overline{\int_{\mathbb{C}}K_{z}(u)S^{\prime}K_{u}(z)d\lambda_{\alpha}(z)}$
$\displaystyle=\overline{S^{\prime}K_{u}(u)}$
$\displaystyle=SK_{u}(u)$
$\displaystyle=e^{\alpha|u|^{2}}\widetilde{S}(u).$
@@
$Tr_{F_{\alpha}^{q}}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(u)\widetilde{S}(u)dA(u).$
∎ We can now prove
@@
Let $1<p,q<\infty$, suppose $\varphi$ is a bounded, compactly supported function on $\mathbb{C}$. Then $T_{\varphi}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$. Furthermore, for any $S\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$ we have:
$Tr_{F_{\alpha}^{q}}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\widetilde{S}(z)dA(z).$ (4.2)
$T_{\varphi}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ follows from Corollary 3.6. Applying the pointwise estimates on the Fock space, we have
@@
$\displaystyle\int_{\mathbb{C}}|\varphi(u)|d\lambda_{\alpha}(u)\int_{\mathbb{C}}|K_{u}(z)||SK_{z}(u)|d\lambda_{\alpha}(z)$
$\displaystyle\lesssim\int_{\mathbb{C}}|\varphi(u)|e^{\frac{\alpha}{2}|u|^{2}}d\lambda_{\alpha}(u)\int_{\mathbb{C}}|K_{u}(z)|e^{-\frac{\alpha}{2}|z|^{2}}dA(z)$
$\displaystyle\simeq\int_{\mathbb{C}}|\varphi(u)|dA(u)$
$\displaystyle<\infty.$
Then the conclusion follows form Theorem 4.3.
The theorem below extends a result in [3, Theorem 9]. The cited theorem asserts that the collection of all Toeplitz operators $T_{\varphi}$ with compactly supported continuous symbols is both trace-norm dense in the trace ideal of $F_{\alpha}^{2}$ and operator-norm dense in $\mathcal{K}(F_{\alpha}^{2})$.
Theorem 4.5.
Let $\mathcal{C}$ denote the set of all Toeplitz operators $T_{\varphi}$ with continuous, compactly supported symbols $\varphi$ on $\mathbb{C}$, and let $1<p,q<\infty$. Then:
  1. (1)
    $\mathcal{C}$ is dense in $\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ with respect to the nuclear norm.
  2. (2)
    $\mathcal{C}$ is norm-dense in $\mathcal{K}(F_{\alpha}^{p},F_{\alpha}^{q})$.
We first prove (2). Assume, for contradiction, that $\mathcal{C}$ is not norm-dense in $\mathcal{K}(F_{\alpha}^{p},F_{\alpha}^{q})$. An application of the Hahn–Banach theorem combined with Corollary 2.5 implies the existence of a nonzero operator $S\in\mathcal{N}(F_{\alpha}^{q},F_{\alpha}^{p})$ satisfying
$\operatorname{Tr}(T_{\varphi}S)=0\qquad\text{for all }T_{\varphi}\in\mathcal{C}.$
@@
$\operatorname{Tr}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\,\widetilde{S}(z)\,dA(z).$
@@
The proof of (1) follows a similar argument. ∎
Funding
Y. Lu was supported by NNSFC (Grant No.12031002), C. Zu was supported by NNSFC (Grant No.12401151), and the Postdoctoral Researcher Foundation of China (Grant No.GZB20240100).
Data availability statement
No data were collected, generated or consulted in connection with this research.
Declarations
Conflict of interest
The authors have no conflict of interest.

### v2

@@ 3 Nuclear Toeplitz operator
Let $1\leq p<\infty$, and let $g$ be an entire function. Then
@@
Then we have $\varphi_{g_{n}}\in(F_{\alpha}^{p})^{\prime}$. Fix $f\in F_{\alpha}^{p}$, since $|g_{n}(z)|\leq|g(z)|$, we can use the dominated convergence theorem to obtain
@@
Hence $\sup_{n}|\varphi_{g_{n}}(f)|<\infty$. It follows from the Principle of Uniform Boundedness that we have
@@
for all $f\in F_{\alpha}^{p}$. It implies that $\varphi_{g}\in(F_{\alpha}^{p})^{\prime}$. Since $(F_{\alpha}^{p})^{\prime}$ can be identified with $F_{\alpha}^{p^{\prime}}$ under the integral pairing $\langle\cdot,\cdot\rangle_{F_{\alpha}^{2}}$, we can find $g^{\prime}\in F_{\alpha}^{p^{\prime}}$ such that
@@
Then we can use the dominated convergence theorem to obtain
$\varphi_{g-g^{\prime}}(z^{n})$ $=\int_{\mathbb{C}}z^{n}\overline{(g-g^{\prime})(z)}e^{-\alpha|z|^{2}}dA(z)$
$=lim_{R\rightarrow{\infty}}\int_{|z|<R}z^{n}\overline{(g-g^{\prime})(z)}e^{-\alpha|z|^{2}}dA(z)$
$=lim_{R\rightarrow{\infty}}\overline{a_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z).$
$=\overline{a_{n}}\int_{\mathbb{C}}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
$=0.$
@@
The other direction follows from Hölder’s inequality.
@@
It is well-known that for $1\leq p<q\leq\infty$, we have the (proper) inclusion $F_{\alpha}^{p}\subset F_{\alpha}^{q}$, and the inclusion is proper. Theorem 3.6 tells us: We can find $f\in F_{\alpha}^{p^{\prime}}$ and $g\in F_{\alpha}^{q}$ such that
@@
The membership of $f$ in $F_{\alpha}^{p^{\prime}}$ is verified by computing two sums. For $1<p^{\prime}\leq 2$,
@@
$\sum_{n=1}^{\infty}|a_{n}|^{p}\left(\frac{n!}{\alpha^{n}}\right)^{\frac{p}{2}}n^{\frac{p}{4}-\frac{1}{2}}=\sum_{k=1}^{\infty}b^{pk}n_{k}^{\frac{p}{4}-\frac{1}{2}+\frac{p}{4}-\frac{p}{2p^{\prime}}}\epsilon_{n_{k}}^{p}\leq\sum_{k=1}^{\infty}\left(\frac{b^{p}}{2^{p}}\right)^{k}<\infty.$
The convergence of these series, together with Theorems 1 and 4 in [MR2215157], implies that $f\in F_{\alpha}^{p^{\prime}}$. Since $a_{n}\neq o\left(\sqrt{\frac{\alpha^{n}}{n!}}n^{\frac{1}{4}-\frac{1}{2q^{\prime}}}\right)\quad(n\rightarrow\infty),$ it follows from Corollary 5 of [MR2215157] that $f\in F_{\alpha}^{p^{\prime}}\setminus F_{\alpha}^{q^{\prime}}$.
@@
$b_{m}=\begin{cases}b^{k}\sqrt{\dfrac{\alpha^{n_{k}}}{(n_{k})!}}\,n_{k}^{\frac{1}{4}-\frac{1}{2q}}\,\epsilon_{n_{k}},&\text{if }m=n_{k},\\ 0,&\text{otherwise,}\end{cases}$
@@
If $\int_{\mathbb{C}}|f(z)g(z)|e^{-\alpha|z|^{2}}\mathrm{d}A(z)<\infty,$ then by the dominated convergence theorem, we obtain
$\int_{\mathbb{C}}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}\mathrm{d}A(z)$ $=\lim_{R\rightarrow\infty}\int_{|z|<R}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}\mathrm{d}A(z)$
$=\lim_{R\rightarrow\infty}\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}a_{n}\overline{b_{m}}\int_{|z|<R}z^{n}\overline{z}^{m}e^{-\alpha|z|^{2}}\mathrm{d}A(z)$
$=\lim_{R\rightarrow\infty}\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}\mathrm{d}A(z).$
From the construction above, we have $a_{n}\overline{b_{n}}\geq 0$ and $\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}\mathrm{d}A(z)$ is monotonically increasing in $R$. The monotone convergence theorem yields:
$\int_{\mathbb{C}}f(z)\overline{g(z)}e^{\alpha|z|^{2}}dA(z)$ $=lim_{R\rightarrow{\infty}}\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
$=\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{\mathbb{C}}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
$=\frac{\pi}{\alpha}\sum_{k=1}^{\infty}\gamma^{2k}\delta_{n_{k}}$
$\geq\frac{\pi}{\alpha}\sum_{k=1}^{\infty}(\frac{\gamma^{2}}{3})^{k}$
$=\infty.$
@@
Finally, we conclude this section with an open problem.
Question 3.8.
@@
4 Trace formula and density theorem
@@
$\mathrm{Tr}(T)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T}(z)dA(z).$
Proof 9.
Let $T=\sum_{j=1}^{\infty}f_{j}\otimes g_{j}$ be a nuclear representation of $T$, where $\left(g_{j}\right)\subset F_{\alpha}^{p},\left(f_{j}\right)\subset F_{\alpha}^{p^{\prime}}$. Then we have $\mathrm{Tr}(T)=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle_{F_{\alpha}^{2}}$. And the trace is independent of the choice of representation. Then we have:
$\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T}(z)dA(z)$ $=\frac{\alpha}{\pi}\int_{\mathbb{C}}\langle Tk_{z},k_{z}\rangle_{F_{\alpha}^{2}}dA(z)$
$=\frac{\alpha}{\pi}\int_{\mathbb{C}}\sum_{j=1}^{\infty}\langle K_{z},f_{j}\rangle_{F_{\alpha}^{2}}\langle g_{j},K_{z}\rangle_{F_{\alpha}^{2}}e^{-\alpha|z|^{2}}dA(z)$
$=\sum_{j=1}^{\infty}\int_{\mathbb{C}}\overline{f_{j}(z)}g_{j}(z)d\lambda_{\alpha}(z)$
$=\mathrm{Tr}(T).$
Notice that the case $p=q=2$ of the following theorem appears in [MR1277446, Theorem 6], but we adopt a different approach here.
@@
$\mathrm{Tr}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\widetilde{S}(z)dA(z).$
Proof 10.
Let $f\in F_{\alpha}^{p}$, we can find a sequence $f_{n}\in K$ such that $f_{n}$ converges to $f$ with $\|\|_{\alpha,p}$. The ser $K$ is defined as (2.2). By assumption (2), we have $T_{\varphi}f_{n}\stackrel{{\scriptstyle\|\cdot\|_{\alpha,q}}}{{\longrightarrow}}T_{\varphi}f$, hence $T_{\varphi}f_{n}(z)\rightarrow T_{\varphi}f(z)$ for every $z\in\mathbb{C}$.
@@
$|\int_{\mathbb{C}}(f_{n}(w)-f(w))\varphi(w)K(z,w)d\lambda_{\alpha}(w)|$
$\leq\frac{\alpha}{\pi}\int_{\mathbb{C}}|(f_{n}(w)-f(w))e^{-\frac{\alpha}{2}|z|^{2}}||\varphi(w)K(z,w)e^{-\frac{\alpha}{2}|z|^{2}}|dA(w)$
$\leq C\|f_{n}-f\|_{\alpha,p}\|\varphi K_{z}\|_{\alpha,p^{\prime}}\rightarrow 0(n\rightarrow\infty).$
@@
$\mathrm{Tr}(T_{\varphi}S)$ $=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T_{\varphi}S}(z)dA(z)$ (4.1)
@@
$\int_{\mathbb{C}}K_{u}(z)SK_{z}(u)d\lambda_{\alpha}(z)$ $=\int_{\mathbb{C}}K_{u}(z)\langle SK_{z},K_{u}\rangle_{F_{\alpha}^{2}}d\lambda_{\alpha}(z)$
$=\int_{\mathbb{C}}K_{u}(z)\langle K_{z},S^{\prime}K_{u}\rangle_{F_{\alpha}^{2}}d\lambda_{\alpha}(z)$
$=\overline{\int_{\mathbb{C}}K_{z}(u)S^{\prime}K_{u}(z)d\lambda_{\alpha}(z)}$
$=\overline{S^{\prime}K_{u}(u)}$
$=e^{\alpha|u|^{2}}\widetilde{S}(u).$
@@
$\mathrm{Tr}_{F_{\alpha}^{q}}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(u)\widetilde{S}(u)dA(u).$
We can now prove
@@
Let $1<p,q<\infty$, and let $\varphi$ be a bounded, compactly supported function on $\mathbb{C}$. Then $T_{\varphi}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$. Furthermore, for any $S\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$, we have:
$\mathrm{Tr}_{F_{\alpha}^{q}}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\widetilde{S}(z)dA(z).$ (4.2)
Proof 11.
$T_{\varphi}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ follows from Theorem 3.4. Applying the pointwise estimates on the Fock space, we have
@@
$\int_{\mathbb{C}}|\varphi(u)|d\lambda_{\alpha}(u)\int_{\mathbb{C}}|K_{u}(z)||SK_{z}(u)|d\lambda_{\alpha}(z)$
$\leq C\int_{\mathbb{C}}|\varphi(u)|e^{\frac{\alpha}{2}|u|^{2}}d\lambda_{\alpha}(u)\int_{\mathbb{C}}|K_{u}(z)|e^{-\frac{\alpha}{2}|z|^{2}}dA(z)$
$=C\int_{\mathbb{C}}|\varphi(u)|dA(u)$
$<\infty.$
Then the conclusion follows from Theorem 4.2.
We now present the proof of Theorem 1.2:
Proof 12 (Proof of Theorem 1.2).
We first prove (2). Assume, for contradiction, that $C$ is not norm-dense in $\mathcal{K}(F_{\alpha}^{p},F_{\alpha}^{q})$. An application of the Hahn-Banach theorem combined with Proposition 2.5 implies the existence of a nonzero operator $S\in\mathcal{N}(F_{\alpha}^{q},F_{\alpha}^{p})$ satisfying
$\mathrm{Tr}(T_{\varphi}S)=0\qquad\text{for all }T_{\varphi}\in C.$
@@
$\mathrm{Tr}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\,\widetilde{S}(z)\,dA(z).$
@@
The proof of (1) follows by a similar argument.